A gas is compressed from 5 L to 3 L under an average pressure of 200000 Pa. During the compression 0.3 kJ of heat is given off. The internal energy of the gas
1
Expert's answer
2020-11-09T10:13:47-0500

Solution

For an isobaric process workdone

W = P(V2 - V1).

"W=200000\times(0.003-0.005)"

"W=-400J"

"\Delta Q=300J"

Energy conservation: ΔU = ΔQ - ΔW.

ΔU = 300-(-400) =700J